I am trying to find the Y so that the line passing through the two points has the given slope. I am not exactly clear on how to find the Y in this problem.
(2,y), (3,3), m=2
This is what I've tried so far
3-y/3-2=2y/1 This is as far as I got. Please help. Thanks

Answer by xcentaur(357)

(2,y), (3,3), m=2


Now suppose we have a line through these two points
slope of the line would be:
M=difference in y-coordinates/difference in x-coordinates
M=(y-3)/(2-3)
M=(y-3)/-1
M=-(y-3)
M=-y+3
M=3-y


But we have been told that the slope is 2
so we put this value in place of M


2=3-y
2-3=-y
-1=-y
1=y
y=1


Hope this helps,
-xC